The Bem Sex Role Inventory (BSRI) provides independent assessments of masculinity and femininity in terms of the respondent's self-reported possession of socially desirable, stereotypically masculine and feminine personality characteristics Alison Konrad and Claudia Harris sought to compare northern U.S. and southern U.S. women on their judgments of the desirability of 40 masculine, feminine, or androgynous traits. Suppose that the following are the scores from a hypothetical sample of northern U.S. women for the attribute Sensitive 3 1 1 23 Calculate the mean, degrees of freedom, variance, and standard deviation for this sample. M E 2.00 df 4
Suppose that the following are the scores from a hypothetical sample of southern U.S. women for the attribute Independent. Calculate the mean, degrees of freedom, variance, and standard deviation for this sample ME 4.50 df s2
Answer #1

1) Sample Mean M : 2

Solution: 10 > x224 Mean-=- 3+112+3 10 2

Degrees of freedom = n-1 = 5-1 = 4

Sample Variance and Sample SD:

Sample Variance s2- (10)2 24-20 ΙΣΧ) び. Sample Standard deviation S-_ (10) 24 - 24 -20

2) Sample Mean: 4.5

Solution: 4 x = 27 2x2 = 133 Mean-=-n 4+7+3+5+3+5 =4.5

Sample Variance and Sample SD:

222. r Sample Variance s2- (27)2 133-6 133-121.5 11.5 Sample Standard deviation S-11 (27) 133 133 121.5 11.5 2.3 = 1.5166